What is the proton (H^+) concentration (C) of a solution of H4P2O7 acid that has a PH of 0.892?

What is the concentration of (P2O7^-4)?
1
Expert's answer
2020-12-01T09:30:24-0500

Using the formula

PH=-log[H^+]

0.892=-log[H^x]

-0.892=log[H^x]

[H^x]= antilog(-0.892)

[H^x]=0.128

To find the conjugate concentration

We will divide the [H^+] conc by 4

Conc of (P2O7^-4)=0.128/4=0.032